# Simultaneous Calibration of Noise Covariance and Kinematics for State Estimation of Legged Robots via Bi-level Optimization

### v1

@@ I Introduction
State estimation is fundamental to the autonomy of robotic systems, providing the basis for planning and control. Robots must infer their states–such as position, velocity, and orientation-from noisy sensor data, often under challenging real-world conditions. Classical approaches such as the Kalman filter and its extensions [1, 2] remain central, and have been successfully deployed in applications ranging from ground vehicles [3] and aerial robots [4] to quadrupeds and humanoids [5, 6, 7, 8]. More recently, factor graph–based methods and smoothing techniques [9, 10] have enabled large-scale estimation for simultaneous localization and mapping (SLAM) [11] and visual–inertial odometry [12, 13].
A persistent challenge in these methods is the specification of noise covariance matrices. The process noise covariance captures uncertainty in the dynamics and actuation, while the measurement noise covariance characterizes sensor error models. In practice, both are difficult to obtain: manufacturer datasheets provide only partial information, while individual sensor calibration and physical system identification are both expensive and task-dependent. As a result, practitioners often rely on manual tuning [7, 14, 15] - a heuristic, time-consuming process that can produce estimators that are either sub-optimal or inconsistent [3] over out-of-distribution tasks. This problem is particularly acute for highly dynamic robots such as quadrupeds and humanoids, whose complex, high-dimensional hybrid dynamics and kinematics, and multi-modal sensor nonlinearities amplify the challenges to obtain precise state estimates [6, 14, 7, 8, 15], especially under kinematic uncertainties.
@@
Several lines of work have been proposed to address this issue. Adaptive filtering methods [16] adjust covariances online based on innovation statistics, while expectation–maximization (EM) frameworks [17, 18, 19] iteratively refine noise parameters to maximize likelihood [20, 21]. Learning sensor parameters or noises parameters [22, 23, 19] is another common approach. However, these methods often either assume no availability of ground truth data or only focus on decoupled calibrations. More importantly, their treatments have not been examined on modern legged robots that have multimodal sensor suites and complex dynamics.
In this work, we propose a bi-level optimization framework for simultaneous covariance and kinematics calibration for robotic state estimation. At the upper level, we treat covariance matrices and uncertain kinematics as optimization variables. At the lower level, we solve a full-information estimator using the Maximum a Posteriori (MAP) formulation given the current covariance estimates and kinematics information. By differentiating through the estimator [24, 25], we directly optimize parameters to be calibrated with respect to trajectory-level objectives of minimizing the errors between the estimates and the measurable ground truth. This approach eliminates manual tuning, enforces physical consistency through structured parameterization and constraints, and generalizes across sensor modalities and robots.
@@ II Related Work
Legged Robotic State Estimation: State estimation for legged robots has been extensively studied due to the challenges posed by intermittent ground contacts, actuator uncertainties, and highly dynamic motions. Early approaches utilize Extended Kalman Filters (EKFs) to fuse proprioceptive sensing - Inertial Measurement Unit (IMU), contact sensors, and joint encoders - with kinematic constraints to achieve drift-reduced estimates on quadrupedal [5], bipedal [26] and humanoid robots [6]. Recent focuses of using EKFs on legged robots employ the invariant properties of the SE(3) manifold, and thus yield more accurate estimates [7, 14]. Another line of work leveraged factor graphs [10, 27] and smoothing or windowed formulations such as Moving Horizon Estimation (MHE) [15] to integrate proprioceptive and exteroceptive sensing data at scale. Despite these advances, most approaches rely on fixed or manually tuned noise covariances, limiting the accuracy of estimates, especially in the presence of kinematic uncertainty. Our work addresses this gap by introducing a bi-level optimization framework that calibrates covariance matrices and uncertain kinematics simultaneously in a principled manner.
@@
Sensor Calibration: Sensor calibration and noise characterization have also been studied extensively in robotics, since the accuracy of estimation strongly depends on the quality of sensor models. Traditional approaches focus on extrinsic and intrinsic calibration of individual sensors, such as cameras [28], LiDARs [29], and IMUs [30]. At the theoretical level, several works address the identification of noise statistics. For instance, expectation–maximization (EM) algorithms iteratively refine noise models to maximize likelihood [17, 18, 19]. Recent efforts explored data-driven estimation of uncertainty, including machine learning to optimize covariances from raw sensory inputs [23, 31]. Despite these advances, most existing techniques commonly are tailored to specific sensors rather than complex legged robots; they often assume restrictive statistical models, or require careful initialization. In contrast, our bi-level optimization jointly calibrates both process and measurement noise with uncertain kinematics across diverse and complex legged robotic platforms.
@@
System Identification: System identification (SysID) has long been a cornerstone of robotics, enabling accurate modeling of dynamics, kinematics, and actuation. Classical techniques for parameter estimation of rigid-body kinematics or dynamics exploit regressions of the geometric or Lagrangian models [32, 33, 34] subject to physical constraints. Gaussian processes or neural networks approximate residuals not captured by physics-based models [35] are commonly used in calibration. Recent works [36, 37] focus on online kinematics calibration via state estimators. Despite significant progress, SysID typically treats noise statistics, model parameters, and estimator design separately. Our bi-level optimization approach bridges this gap by jointly calibrating positive definite covariance matrices and kinematic parameters, highlighting the natural synergy between system identification and state estimation.
@@ III Preliminaries
We first introduce the models of legged state estimation that are widely used in the literature [6, 5, 14, 15, 7]. Then, we will elaborate the preliminaries of the mathematics on formulating state estimation as an optimization problem that we will employ in our calibration work. Throughout, $\mathrm{W}$ denotes the world frame and $\mathrm{B}$ the body (base) frame.
@@ III-A2 Measurement Model
where $\cdot$$_{\times}$ denotes skew-symmetric map, $\delta_{\mathrm{pf}}\!\sim\!\mathcal{N}(0,R_{\mathrm{pf}})$ and $\delta_{\mathrm{vf}}\!\sim\!\mathcal{N}(0,R_{\mathrm{vf}})$ encapsulates kinematics errors and encoder/gyroscope noises [26]. Here $\tilde{(\cdot)}$ denotes a raw sensor measurement.
@@ III-B State Estimation as an MAP Problem
The major drawback of FIE is that its computational cost grows with the horizon $T$, making online use practically intractable. MHE [15] instead uses a fixed past window of measurements and repeatedly solves the problem on a shifting horizon (as in MPC), yielding bounded and predictable cost. Because legged estimation is real-time, EKF variants with analytical updates are commonly used instead of windowed optimization. For offline calibration, we exploit all available data and therefore adopt FIE.
@@ IV Problem Statement
With properly selections of the noise covariance matrices $Q$ and $R$, the commonly applied estimators such as Extended Kalman Filters (EKFs) [5, 6], invariance EKFs [14, 7], and MHE [15] can yield satisfying estimates. However, it is widely known that extensive expert tuning of $Q$ and $R$ is inevitable on complex robots, and it can easily lead to sub-optimal, inconsistent or even diverging estimates on real robots. Additionally, inaccurate kinematics can further degrade estimate stability and precision. Therefore, we aim to provide an automatic and tuning-free method to optimally calibrate $Q$ and $R$ along with uncertain kinematics so that any reasonable choices of the estimators, regardless of EKFs or MHEs depending on the computational budget, can yield precise and consistent estimation on diverse set of tasks.
@@ V Method
We now present our method to solve the bi-level optimization problem in (12). We adopt an iterative optimization approach: the lower-level problem (8) is solved using a standard interior-point method [38], while the upper-level problem leverages its solution and Karush–Kuhn–Tucker (KKT) condition to update the parameters $\bm{\theta}$ via an adaptive Frank–Wolfe algorithm that iteratively refines the estimate until convergence. The index $t\in\mathbb{N}$ denotes the Frank–Wolfe iteration number.
@@ VI Results
We evaluate our calibration approach on three robots: a bipedal robot STRIDE [15] and two quadrupedal robot Go1 and B1 from Unitree. The data used from STRIDE and Go1 are obtained from simulation in Matlab and MuJoCo environment in their open source repository [15]. We obtained the data of B1 on hardware in the motion capture room with 12 Opti-track Cameras, with a combination of Prime 13 and 22.
The bi-level optimization is implemented in Python and C++. The outer loop iteration as LMO is solved by MOSEK; the inner loop FIE is solved with IPOPT [38]. The analytical gradients are generated via CasADi [39] and Pinocchio [40].

### v2

@@ I Introduction
State estimation is fundamental to the autonomy of robotic systems, providing the basis for planning and control. Robots must infer their states–such as position, velocity, and orientation-from noisy sensor data, often under challenging real-world conditions. Classical approaches such as the Kalman filter and its extensions [kalman1960, bai2023state] remain central, and have been successfully deployed in applications ranging from ground vehicles [thrun2005probabilistic] and aerial robots [mohamed2021momentum] to quadrupeds and humanoids [bloesch2013state, rotella2014state, he2025invariant, HLIP]. More recently, factor graph–based methods and smoothing techniques [kaess2012isam2, dellaert2017factor] have enabled large-scale estimation for simultaneous localization and mapping (SLAM) [cadena2017past] and visual–inertial odometry [mur2017orb, usenko2019visual].
A persistent challenge in these methods is the specification of noise covariance matrices. The process noise covariance captures uncertainty in the dynamics and actuation, while the measurement noise covariance characterizes sensor error models. In practice, both are difficult to obtain: manufacturer datasheets provide only partial information, while individual sensor calibration and physical system identification are both expensive and task-dependent. As a result, practitioners often rely on manual tuning [he2025invariant, hartley2020contact, kang24] - a heuristic, time-consuming process that can produce estimators that are either sub-optimal or inconsistent [thrun2005probabilistic] over out-of-distribution tasks. This problem is particularly acute for highly dynamic robots such as quadrupeds and humanoids, whose complex, high-dimensional hybrid dynamics and kinematics, and multi-modal sensor nonlinearities amplify the challenges to obtain precise state estimates [rotella2014state, hartley2020contact, he2025invariant, HLIP, kang24], especially under kinematic uncertainties.
@@
Several lines of work have been proposed to address this issue. Adaptive filtering methods [9044358] adjust covariances online based on innovation statistics, while expectation–maximization (EM) frameworks [em_mhe_2017, 10472604, yap2008simultaneous] iteratively refine noise parameters to maximize likelihood [6630924, 6482657]. Learning sensor parameters or noises parameters [wong2020variational, qadri2024learning, yap2008simultaneous] is another common approach. However, these methods often either assume no availability of ground truth data or only focus on decoupled calibrations. More importantly, their treatments have not been examined on modern legged robots that have multimodal sensor suites and complex dynamics.
In this work, we propose a bi-level optimization framework for simultaneous covariance and kinematics calibration for robotic state estimation. At the upper level, we treat covariance matrices and uncertain kinematics as optimization variables. At the lower level, we solve a full-information estimator using the Maximum a Posteriori (MAP) formulation given the current covariance estimates and kinematics information. By differentiating through the estimator [amos2017optnet, icml2022implicit], we directly optimize parameters to be calibrated with respect to trajectory-level objectives of minimizing the errors between the estimates and the measurable ground truth. This approach eliminates manual tuning, enforces physical consistency through structured parameterization and constraints, and generalizes across sensor modalities and robots.
@@ II Related Work
Legged Robotic State Estimation: State estimation for legged robots has been extensively studied due to the challenges posed by intermittent ground contacts, actuator uncertainties, and highly dynamic motions. Early approaches utilize Extended Kalman Filters (EKFs) to fuse proprioceptive sensing - Inertial Measurement Unit (IMU), contact sensors, and joint encoders - with kinematic constraints to achieve drift-reduced estimates on quadrupedal [bloesch2013state], bipedal [teng2021legged] and humanoid robots [rotella2014state]. Recent focuses of using EKFs on legged robots employ the invariant properties of the SE(3) manifold, and thus yield more accurate estimates [he2025invariant, hartley2020contact]. Another line of work leveraged factor graphs [dellaert2017factor, yang2024real] and smoothing or windowed formulations such as Moving Horizon Estimation (MHE) [kang24] to integrate proprioceptive and exteroceptive sensing data at scale. Despite these advances, most approaches rely on fixed or manually tuned noise covariances, limiting the accuracy of estimates, especially in the presence of kinematic uncertainty. Our work addresses this gap by introducing a bi-level optimization framework that calibrates covariance matrices and uncertain kinematics simultaneously in a principled manner.
@@
Sensor Calibration: Sensor calibration and noise characterization have also been studied extensively in robotics, since the accuracy of estimation strongly depends on the quality of sensor models. Traditional approaches focus on extrinsic and intrinsic calibration of individual sensors, such as cameras [rehder2016extending], LiDARs [li2023automatic], and IMUs [fu2021high]. At the theoretical level, several works address the identification of noise statistics. For instance, expectation–maximization (EM) algorithms iteratively refine noise models to maximize likelihood [em_mhe_2017, 10472604, yap2008simultaneous]. Recent efforts explored data-driven estimation of uncertainty, including machine learning to optimize covariances from raw sensory inputs [qadri2024learning, wang2023neural]. Despite these advances, most existing techniques commonly are tailored to specific sensors rather than complex legged robots; they often assume restrictive statistical models, or require careful initialization. In contrast, our bi-level optimization jointly calibrates both process and measurement noise with uncertain kinematics across diverse and complex legged robotic platforms.
@@
System Identification: System identification (SysID) has long been a cornerstone of robotics, enabling accurate modeling of dynamics, kinematics, and actuation. Classical techniques for parameter estimation of rigid-body kinematics or dynamics exploit regressions of the geometric or Lagrangian models [leboutet2021inertial, 7987066, khorshidi2025physically] subject to physical constraints. Gaussian processes or neural networks approximate residuals not captured by physics-based models [hwangbo2019learning] are commonly used in calibration. Recent works [yang2022online, burgul2024online] focus on online kinematics calibration via state estimators. Despite significant progress, SysID typically treats noise statistics, model parameters, and estimator design separately. Our bi-level optimization approach bridges this gap by jointly calibrating positive definite covariance matrices and kinematic parameters, highlighting the natural synergy between system identification and state estimation.
@@ III Preliminaries
We first introduce the models of legged state estimation that are widely used in the literature [rotella2014state, bloesch2013state, hartley2020contact, kang24, he2025invariant]. Then, we will elaborate the preliminaries of the mathematics on formulating state estimation as an optimization problem that we will employ in our calibration work. Throughout, $\mathrm{W}$ denotes the world frame and $\mathrm{B}$ the body (base) frame.
@@ III-A2 Measurement Model
where $\cdot$$_{\times}$ denotes skew-symmetric map, $\delta_{\mathrm{pf}}\!\sim\!\mathcal{N}(0,R_{\mathrm{pf}})$ and $\delta_{\mathrm{vf}}\!\sim\!\mathcal{N}(0,R_{\mathrm{vf}})$ encapsulates kinematics errors and encoder/gyroscope noises [teng2021legged]. Here $\tilde{(\cdot)}$ denotes a raw sensor measurement.
@@ III-B State Estimation as an MAP Problem
The major drawback of FIE is that its computational cost grows with the horizon $T$, making online use practically intractable. MHE [kang24] instead uses a fixed past window of measurements and repeatedly solves the problem on a shifting horizon (as in MPC), yielding bounded and predictable cost. Because legged estimation is real-time, EKF variants with analytical updates are commonly used instead of windowed optimization. For offline calibration, we exploit all available data and therefore adopt FIE.
@@ IV Problem Statement
With properly selections of the noise covariance matrices $Q$ and $R$, the commonly applied estimators such as Extended Kalman Filters (EKFs) [bloesch2013state, rotella2014state], invariance EKFs [hartley2020contact, he2025invariant], and MHE [kang24] can yield satisfying estimates. However, it is widely known that extensive expert tuning of $Q$ and $R$ is inevitable on complex robots, and it can easily lead to sub-optimal, inconsistent or even diverging estimates on real robots. Additionally, inaccurate kinematics can further degrade estimate stability and precision. Therefore, we aim to provide an automatic and tuning-free method to optimally calibrate $Q$ and $R$ along with uncertain kinematics so that any reasonable choices of the estimators, regardless of EKFs or MHEs depending on the computational budget, can yield precise and consistent estimation on diverse set of tasks.
@@ V Method
We now present our method to solve the bi-level optimization problem in (12). We adopt an iterative optimization approach: the lower-level problem (8) is solved using a standard interior-point method [wachter2006implementation], while the upper-level problem leverages its solution and Karush–Kuhn–Tucker (KKT) condition to update the parameters $\bm{\theta}$ via an adaptive Frank–Wolfe algorithm that iteratively refines the estimate until convergence. The index $t\in\mathbb{N}$ denotes the Frank–Wolfe iteration number.
@@ VI Results
We evaluate our calibration approach on three robots: a bipedal robot STRIDE [kang24] and two quadrupedal robot Go1 and B1 from Unitree. The data used from STRIDE and Go1 are obtained from simulation in Matlab and MuJoCo environment in their open source repository [kang24]. We obtained the data of B1 on hardware in the motion capture room with 12 Opti-track Cameras, with a combination of Prime 13 and 22.
The bi-level optimization is implemented in Python and C++. The outer loop iteration as LMO is solved by MOSEK; the inner loop FIE is solved with IPOPT [wachter2006implementation]. The analytical gradients are generated via CasADi [Andersson2019] and Pinocchio [carpentier:hal-01866228].